OpenStudy (anonymous):

Suppose the sales ofa particular brand of appliance satify the linear model y = 80x + 300, where y represents the number of sales in year x, with x + 0 corresponding to 1982, with x +1 corresponding to 1983, etc. Find the number of sales in 1990. A. 1800 B. 860 C. 940 D. 1880

OpenStudy (anonymous):

y = 80x + 300 when x=8 (1990) y=80*8+300 =640+300 =940 C.

OpenStudy (anonymous):

thank you

OpenStudy (king):

vince dont delete this post now..pls..

OpenStudy (anonymous):

ok. just curious why

OpenStudy (king):

why do u wanna delete it?

OpenStudy (anonymous):

Sorry, I was just curious

OpenStudy (anonymous):

If you leave this post here, other people with similar questions can view this as an example to help solve their problems. :)

OpenStudy (anonymous):

Oh, ok. I'm sorry

OpenStudy (king):

exactly!

OpenStudy (anonymous):

No worries, it's not a problem :D

OpenStudy (anonymous):

cool

OpenStudy (king):

and if u are scared that some one will make fun of u.....dont be scared!

OpenStudy (anonymous):

oh i'm not :)